Three dices are rolled once what is the probability that the product of the faces is more than 215
Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

the only way to get a product greater than 215 is if each die has a value of 6 when it is rolle.
in that case, the product would be equal to 6^3 = 216.
if only one of the die had something less than 6, like 5, then the product woudl be equal to 6^2 * 5 = 180.
this is decidedly less than 215.
the probability of getting all sixes would be equal to 1/6 * 1/6 * 1/6 = 1/216.

Answer by ikleyn(52795)

.
Three dices are rolled once what is the probability that the product of the faces is more than 215
~~~~~~~~~~~~~


For your info


            Use  " die "  as single;

            Use  " dice "  as plural.



Do not use  " dices "  as plural.   It is incorrect.